In the case of a game of chance, every 20th person usually wins an amount of more than EUR10000. How many players have to compete so that with a probability of 95% at least one more than ¨ EUR 10,000 wins?

In a large sports competition, the participants are checked for doping. Experience has shown that 4 out of 75 participants consume illicit substances. A test should check whether the athlete ¨
fights fairly. It is known that this test has a probability of 0.97 to convince a doped athlete as such. However, 2% of athletes are also wrongly accused of doping by the test. What is the probability that a positive athlete actually doped?
Answer #1

proportion of doped participants – 1/5 PD) proportion of otoped Plaad= 1- P10)=1-4 / 4 = 2/15 The test result of test P ( pos